OpenStudy (anonymous):

Bill can hammer 20 nails in 6 minutes. Jeff can do the same job in only 5 minutes . How long will it take them to finish if Bill hammers the first 5 nails, then Jeff hammers for 3 minutes then Bill finishes the job? A. 4.6 minutes B. 5.0 minutes C. 5.4 minutes D. 5.8 minutes E. 6.0 minutes

OpenStudy (perl):

Bills rate = 20 nails / 6 minutes Jeff's rate = 20 nails / 5 minutes

OpenStudy (perl):

now we can use the formula distance = rate * time, (notice that it doesn't have to be distance, it can be number of nails )

OpenStudy (anonymous):

i gt 3.333 n 4!!

OpenStudy (abhisar):

Hello @dinisha !

OpenStudy (abhisar):

First of all lets calculate how much time bill needs to hammer one nail. Can u do that for me pls ?

OpenStudy (abhisar):

5.4 mins will be the answer

OpenStudy (abhisar):

U there @dinisha ?

OpenStudy (anonymous):

hei ya @Abhisar but i gt 5.8 minutes!!

OpenStudy (abhisar):

okay, lets c, maybe i am wrong....

OpenStudy (abhisar):

First u need to calculate how much time bill needs to hammer one nail.

OpenStudy (abhisar):

Its 20/6 = 0.3 mins

OpenStudy (abhisar):

agree ?

OpenStudy (abhisar):

Now in 3 minutes jeff will hammer 12 nails, so 3 nails more left

OpenStudy (abhisar):

For hammering these 3 nails, bill will take 6/20 X 3 = 0.9 mins

OpenStudy (abhisar):

Now total time = 0.9 + 3 +1.5 = 5.4

OpenStudy (anonymous):

ohh oky ya2 nw i understand.. i went wrong at the part or 0.3 minutes.. thank u @Abhisar

OpenStudy (abhisar):

Did you know you can give medals to people who help you with questions? Just click the \(\color{blue}{\text{blue "best response" button}}\) to show appreciation for the people who help you.

OpenStudy (abhisar):

\(\Huge\text{Anytime !}\) \(\huge\ddot\smile\)